Can you help me to solve these two equations by using the substitution method.
2x + 4y = -6
x - 3y = 7

x = 2y + 2
2x + 3y = 11

Answer by mananth(16946)

2 x + 4 y = -6 .............1
4 y = -6 + -2 x
/ 4 =
y = -1.5 + -0.5 x
1 x + -3 y = 7 .....................2
Substitute y in (2)
1 x + -3 ( -1.5 + -0.5 x ) = 7
1 x + 4.5 + 1.5 x = 7
1 x 1.5 x = 2.5
2.5 x = 2.5
/ 2.5
x= 1
Plug the value of x in (1)
2 x + 4 y = -6
2 * 1 + 4 y = -6
2 + 4 y = -6
4 y = -8
/ 4
y= -2
x=1,y=-2



----------------



m.ananth@hotmail.ca